### v3

# Parameter estimation of range-migrating targets using OTFS signals from LEO satellites

Abstract
This study investigates a communication-centric integrated sensing and communication system that utilizes orthogonal time-frequency space (OTFS) modulated signals emitted by low Earth orbit satellites to estimate the parameters of space targets experiencing range migration, hereinafter referred to as high-speed targets. Leveraging the signal samples produced by off-the-shelf OTFS demodulators, we derive a novel input-output model for the echo generated by a high-speed target when ideal and rectangular shaping filters are employed. Our findings reveal that the target response exhibits a sparse structure in the delay-Doppler domain, whose support is determined by the target initial-range and range-rate. Range migration induces a structured spread of this response, which is explicitly characterized in the paper and differs from that in previous models. We propose an approximate implementation of the maximum likelihood estimator for the target initial-range, range-rate, and amplitude. The estimation process first obtains coarse information on the target response using a block orthogonal matching pursuit algorithm, followed by a refinement step based on a bank of matched filters focused on a smaller initial-range/range-rate region. The proposed single-target procedure is extended to multiple targets via iterative estimation, reconstruction, and cancellation of dominant echoes. Finally, numerical examples are provided to evaluate the estimation performance.
I Introduction
Future wireless networks and Internet-of-Things (IoT) systems will utilize non-terrestrial stations to ensure seamless and ubiquitous connectivity [1]. Among space transceivers, low Earth orbit (LEO) satellites are attracting significant attention due to their low latency, high data rates, and capability to support integrated sensing and communication (ISAC) [2, 3, 4], enabling applications such as environmental monitoring, global navigation, and space exploration.
Orthogonal time frequency space (OTFS) modulation is a promising transmission scheme for future LEO constellations [5]. OTFS-based transceivers exploit channel diversity across both time and frequency, thus providing improved robustness in doubly-dispersive fading channels compared to orthogonal frequency division multiplexing [6, 7]. They are also suitable for radar sensing, as the underlying signal processing resembles that of pulsed-Doppler radars [8]; in particular, the range and range-rate parameters of the detected objects are inferred from the estimated delay-Doppler channel response. [^1] In this work, the range is defined as the propagation distance of the signal path, while the range-rate is defined as its time derivative. Hence, the range and range-rate provide a convenient physical parameterization of the delay and Doppler shifts along the propagation path, respectively. These features enable OTFS-based LEO satellites to support ISAC in various forms [9, 10]. Communication-centric ISAC appears particularly promising in this context, as existing signals can be reused for sensing at low cost without altering the underlying communication protocol [11, 12, 13, 14, 15]; this approach is considered in this study.
A large body of work has addressed signal detection and channel estimation for OTFS systems by exploiting the inherent sparsity of the delay-Doppler representation when range migration over the processed frame can be ignored, an assumption that is usually valid for terrestrial scenarios; see [16] and references therein. For example, [6] derived a widely-used input-output signal model under the assumption of integer range and fractional range-rate for both ideal and rectangular shaping filters, along with a message passing algorithm for data detection. [^2] Integer (fractional) range or range-rate means that the range or range-rate is an integer (fractional) multiple of the respective resolution. Based on this model, pilot-aided channel estimation schemes have been proposed in [17], where guard symbols are introduced to avoid interference between pilot and data symbols at the receiver. Sparse Bayesian learning approaches have been developed in [18] and extended in [19] to handle both fractional range and range-rate. On the sensing side, [11] proposed to exploit OTFS communication signals for radar sensing within a communication-centric ISAC framework. This idea has been further developed in several directions, including the derivation of Cramér–Rao lower bounds for range and range-rate estimation [20], the analysis of pulse shaping effects [21], the design of sparse recovery algorithms with fractional refinement [22], and atomic norm-based formulations for target parameter estimation [23]. The 3D structured sparsity of MIMO-OTFS channels has been investigated in [24], where the downlink channel estimation problem is formulated as a sparse signal recovery task. In addition, OTFS signals have been employed for joint localization and communication in [25] and for target tracking in [26].
Assuming the range of each propagation path to be constant may no longer be valid in satellite-based sensing. For instance, in a scenario with the communication transmitter on a LEO satellite and the radar receiver on an airborne platform, non-cooperative space targets (e.g., debris, asteroids, spacecraft, or missiles) may exhibit range rates of several km/s. When considering a system with $M=512$ subcarriers, $N=128$ symbol intervals, a subcarrier spacing of $\Delta=15$ kHz, a symbol interval of $T=1/\Delta$, and a range rate of $\bar{v}=15$ km/s, the range variation over the OTFS frame is $\bar{v}NT=128$ m, which greatly exceeds the range resolution $c/(M\Delta)\approx 39$ m, where $c$ is the speed of light.
This work investigates the use of OTFS communication signals for estimating the parameters of high-speed targets undergoing range migration within the OTFS frame. We adopt a communication-centric approach in which the waveform is not modified to preserve communication performance, and we focus on characterizing the target echo model, designing the radar receiver, and analyzing estimation performance. Sensing high-speed targets is challenging, and prior radar studies have primarily relied on unmodulated or modulated coherent pulse trains (typically, a train of linear frequency-modulated pulses) and developed motion-compensation strategies in the time or frequency domain to coherently integrate consecutive pulses [27, 28, 29, 30]. However, the impact of range migration on OTFS signals remains unexplored, and existing motion-compensation techniques cannot be directly applied in the delay-Doppler domain under standard modulation and demodulation processing. This study aims to provide initial insights into parameter estimation for high-speed targets using OTFS communication signals, without altering the transceiver structure. The major contributions are summarized below.
  • We derive a model for the echo produced by a high-speed target in the delay-Doppler domain, considering both ideal and rectangular shaping filters; the target range is assumed constant over a block of a few symbols, but changes from block to block consistently with the range-rate. Ideal filters yield a circular convolution model, whereas rectangular filters break this convolution structure because of inter-carrier interference (ICI) and inter-symbol interference (ISI). This model subsumes the one in [6] when the target-range migration is negligible.
  • We demonstrate that the response of high-speed targets maintains a sparse structure in the delay-Doppler domain, determined by its initial-range (i.e., the target range at the beginning of the OTFS frame) and range-rate. Specifically, range migration results in an expanded support of this response not captured by existing models.
  • We propose an approximate implementation of the maximum likelihood (ML) estimator for target parameters, which avoids exhaustive search over the entire parameter space. First, a block orthogonal matching pursuit (BOMP) algorithm [31] is used to obtain coarse information on the initial-range and range-rate. Next, a refined estimate of the target parameters is derived using a bank of matched filters focused on a narrower region. The proposed two-step estimator can be combined with the CLEAN algorithm in [32, 33, 34, 35, 12] to handle multiple targets.
  • Numerical simulations are conducted to validate the proposed echo model and assess the performance of the proposed two-step estimator under different system configurations and numbers of targets, also in comparison with the estimators in [17] and [24].
Organization
The remainder of the paper is organized as follows. Sec. II presents the system description. Sec. III presents the delay–Doppler input–output model under ideal and rectangular shaping filters. Sec. IV contains the proposed procedure for target parameter estimation. Sec. V discusses some numerical examples. Sec. VI contains the conclusions. Finally, the appendices contain some mathematical derivations.
Notation
In the following, $\mathbb{R}$ and $\mathbb{C}$ are the set of real and complex numbers, respectively. Column vectors and matrices are denoted by lowercase and uppercase boldface letters, respectively. The symbols $(\,\cdot\,)^{*}$, $(\,\cdot\,)^{\mathsf{T}}$, and $(\,\cdot\,)^{\mathsf{H}}$ denote conjugate, transpose, and conjugate-transpose, respectively. $\bm{I}_{M}$ is the $M\times M$ identity matrix. $(\bm{a})_{i}$ denotes the $i$-th entry of the vector $\bm{a}$. $(\bm{A})_{i,j}$ denotes the entry in the $i$-th row and $j$-th column of the matrix $\bm{A}$. $(\bm{A}_{1}\ \cdots\ \bm{A}_{n})$ and $(\bm{A}_{1};\ \cdots\ ;\bm{A}_{n})$ denote the horizontal and vertical concatenations of the matrices $\bm{A}_{1},\ldots,\bm{A}_{n}$. $\mathbbm{1}_{\mathcal{A}}$ is the indicator function of the condition $\mathcal{A}$, i.e., $\mathbbm{1}_{\mathcal{A}}=1$, if $\mathcal{A}$ holds true, and $\mathbbm{1}_{\mathcal{A}}=0$ otherwise. $\cdot\,]_{M}$ denotes the mod $M$ operation. $\delta(t)$ and $\delta[n]$ denote the Dirac delta function and the discrete unit sample function, respectively. $\Re\left\{\cdot\right\}$ denotes the real part. Finally, $\mathrm{i}$, $\odot$, and $\mathrm{E}[\,\cdot\,]$ denote the imaginary unit, the Schur product, and the statistical expectation, respectively.
II System description
Consider a system consisting of a communication transmitter and a radar receiver. The communication transmitter is mounted on a LEO satellite, uses an OTFS modulation, and operates at a carrier frequency $c/\lambda$, where $\lambda$ is the carrier wavelength; the emitted waveform is specified by the underlying communication protocol and is not modified here. The radar receiver aims to estimate the initial-range, range-rate, and amplitude of high-speed targets illuminated by the communication transmitter; it may be mounted on the same platform hosting the communication transmitter (monostatic configuration) or on a separate device (bistatic configuration), e.g., another LEO satellite, an aircraft, a high altitude platform, or a terrestrial station. The receiver is synchronized with the transmitter and knows the emitted symbols: for a monostatic configuration, this is readily verified [12], whereas, for a bistatic configuration, this requires some coordination [^3] When the radar receiver is on a different LEO satellite, the coordination may be established via an inter-satellite link. or a separate reference channel [36, 37, 38]. For illustration, in this study we consider the bistatic configuration depicted in Fig. 1.
We assume point-like targets, each generating a single line-of-sight propagation path. The point-like model is appropriate when the target extent is small relative to the range resolution and allows us to isolate and analyze the effects of range migration. The single-path model is consistent with the considered space-sensing scenario, where secondary scattering components are typically negligible or absent. Extensions to distributed targets or to more general propagation conditions involving multiple echoes from the same target are left for future work. For clarity of exposition, we initially focus on a single target, while multiple targets are addressed in Sec. IV-B.
Figure 1: System geometry.
II-A Transmit signal
Consider an OTFS frame with bandwidth $M\Delta$ and duration $NT$, where $M$ and $N$ are the number of subcarriers and symbol intervals, respectively, $\Delta$ is the subcarrier spacing, and $T=1/\Delta$ is the symbol interval. Accordingly, the corresponding range and range-rate resolutions are [36]
$R_{\rm r}$ $=c/(M\Delta),$ (1a)
$R_{\rm rr}$ $=\lambda/(NT),$ (1b)
[…]
$x(t)=\sum_{n=0}^{N-1}\underbrace{\sum_{m=0}^{M-1}X_{\rm TF}[n,m]g_{\rm tx}\!\left(t-nT\right)\mathrm{e}^{\mathrm{i}2\pi m\Delta(t-nT)}}_{x_{n}(t)},$ (2)
where $x_{n}(t)$ is the signal emitted in the $n$-th symbol interval, $X_{\rm TF}[n,m]$ is the time-frequency symbol sent in the $n$-th symbol interval on subcarrier $m$, and $g_{\rm tx}(t)$ is the transmit shaping filter with unit energy, effective temporal duration $T$, and effective spectral width $\Delta$. The time-frequency sequence $X_{\rm TF}[n,m]$ is obtained as the inverse Symplectic Finite Fourier Transform of the delay-Doppler sequence $X_{\rm DD}[k,l]$, i.e.,
$X_{\rm TF}[n,m]=\frac{1}{\sqrt{NM}}\sum_{k=0}^{N-1}\sum_{l=0}^{M-1}X_{\rm DD}[k,l]\mathrm{e}^{\mathrm{i}2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)},$ (3)
for $n=0,\ldots,N-1$ and $m=0,\ldots,M-1$. By considering $g_{\rm tx}(t)$ time limited in $[0,T]$, then $x(t)$ has support $[0,NT]$. Also, by assuming the spectrum of $g_{\rm tx}(t)$ is contained in $0,\Delta]$, then the spectrum of $x(t)$ is contained in [^5] Since a time-limited signal cannot be strictly band-limited, some spectral leakage outside $0,M\Delta]$ is present. In practice, the transmit shaping filter may not be exactly zero outside $[0,T]$ to reduce such spectral leakage; in this latter case, $x(t)$ is only approximately contained in $[0,NT]$. $0,M\Delta]$.
II-B Target kinematics and design assumptions
During an OTFS frame, the target, the communication transmitter, and the radar receiver can be assumed to move with a constant velocity. The position of the target at time $t\in[0,NT]$ is $\bm{p}(t)=\bm{p}-\bm{v}t\in\mathbb{R}^{3}$, where $\bm{p}$ and $\bm{v}$ are the initial position and the velocity, respectively; similarly, the positions of communication transmitter and the radar receiver at time $t\in[0,NT]$ are $\bm{p}_{\rm tx}(t)=\bm{p}_{\rm tx}-\bm{v}_{\rm tx}t\in\mathbb{R}^{3}$ and $\bm{p}_{\rm rx}(t)=\bm{p}_{\rm rx}-\bm{v}_{\rm rx}t\in\mathbb{R}^{3}$, respectively. Accordingly, for $t\in[0,NT]$, the target range is
$\bar{r}(t)=\left\|\bm{p}(t)-\bm{p}_{\rm tx}(t)\right\|+\left\|\bm{p}(t)-\bm{p}_{\rm rx}(t)\right\|.$ (4)
Upon ignoring the quadratic and higher order terms in the Taylor expansion of the above expression, we have that [28]
$\bar{r}(t)=\bar{d}-\bar{v}t,$ (5)
where $\bar{d}$ and $\bar{v}$ are the initial-range and the range-rate, respectively, defined as
$\bar{d}$ $=\left\|\bm{p}-\bm{p}_{\rm tx}\right\|+\left\|\bm{p}-\bm{p}_{\rm rx}\right\|,$ (6a)
$\bar{v}$ $=\frac{\left(\bm{p}-\bm{p}_{\rm tx}\right)^{\mathsf{T}}\!\left(\bm{v}_{\rm tx}-\bm{v}\right)}{\left\|\bm{p}-\bm{p}_{\rm tx}\right\|}+\frac{\left(\bm{p}-\bm{p}_{\rm rx}\right)^{\mathsf{T}}\!\left(\bm{v}_{\rm rx}-\bm{v}\right)}{\left\|\bm{p}-\bm{p}_{\rm rx}\right\|}.$ (6b)
Hereafter, we make the following design assumptions.
II-B1 Unambiguous initial-range and range-rate intervals
Denote by $\bar{r}_{\min}$ and $\bar{r}_{\max}$ the minimum and maximum initial-ranges, respectively, and by $\bar{v}_{\min}$ and $\bar{v}_{\max}$ the minimum and maximum range-rates, respectively. These values are tied to the prior uncertainty on the target location and mobility, as resulting from a previous detection phase (if the radar is operating in search mode) or as predicted by the tracking stage (if the radar is operating in tracking mode), and are assumed known. Also, let $r_{\max}=\bar{r}_{\max}-\bar{r}_{\min}$ and $v_{\max}=\bar{v}_{\max}-\bar{v}_{\min}$ be the length of the inspected initial-range and range-rate intervals, respectively. Then, we assume
$\tau_{\max}$ $=\frac{r_{\max}}{c}<T,$ (7a)
$\nu_{\max}$ $=\frac{v_{\max}}{\lambda}<\Delta,$ (7b)
so as to avoid ambiguity in the estimation of the initial-range and the range-rate of the target, respectively. [^6] For the parameters adopted in Sec. V ($\Delta=15$ kHz, $M=512$, $N=128$, carrier frequency $4$ GHz), we have $c/\Delta\approx 19.9$ km and $\lambda\Delta\approx 1.12$ km/s, respectively. By varying $\bar{r}_{\min}$ and $\bar{v}_{\min}$, any initial-range interval of length $r_{\max}<19.9$ km and any range-rate interval of length $v_{\max}<1.12$ km/s can be jointly inspected without ambiguity.
II-B2 Blockwise constant-range approximation
Let $B$ be a divisor of $N$. Then, the OTFS frame can be partitioned into $B$ equal blocks of $N/B$ symbol intervals. For receiver design, the target range is assumed constant within each block. Accordingly, we write
$\bar{r}(t)=\bar{r}(bT_{B}),\quad\forall t\in[bT_{B},(b+1)T_{B}),$ (8)
where $T_{B}=NT/B$ and $b=0,\ldots,B-1$. Increasing $B$ improves the model accuracy, particularly when the target exhibits higher range-rates or when the OTFS frame duration is long (see Sec. V for further discussion). A practical design guideline is to select $B$ such that the range variation over each block, i.e., $\bar{v}_{\max}NT/B$, is significantly smaller than the range resolution $R_{\rm r}$ (see also Sec. V). Prior studies assume $B=1$, which is appropriate only for applications in which target-range migration within the OTFS frame can be neglected.
II-C Received signal
Upon neglecting the noise contribution, the radio-frequency target echo at the radar receiver is
$\bar{y}_{\rm RF}(t)=\Re\left\{\bar{\alpha}x\left(t-\frac{\bar{r}(t)}{c}\right)\mathrm{e}^{\mathrm{i}2\pi\frac{c}{\lambda}\left(t-\frac{\bar{r}(t)}{c}\right)}\right\},$ (9)
where $\bar{\alpha}\in\mathbb{C}$ is the target amplitude (accounting for the radiated energy, the transmit and receive antenna gains, the target radar cross-section, and any signal attenuation). For convenience, we advance the time axis by $\bar{r}_{\min}/c$, thus obtaining
$\bar{y}_{\rm RF}\left(t+\frac{\bar{r}_{\min}}{c}\right)=\Re\left\{y(t)\mathrm{e}^{\mathrm{i}2\pi\frac{c+\bar{v}_{\min}}{\lambda}t}\right\},$ (10)
where
$y(t)=\bar{\alpha}x\left(t-\frac{r(t)}{c}\right)\mathrm{e}^{-\mathrm{i}2\pi\frac{r(t)}{\lambda}}\mathrm{e}^{-\mathrm{i}2\pi\frac{\bar{v}_{\min}}{\lambda}t}$ (11)
is the baseband representation of the target echo with respect to the reference frequency $\frac{c+\bar{v}_{\min}}{\lambda}$ and $r(t)=\bar{r}(t)-\bar{r}_{\min}$ is the excess target range (i.e., the target range in excess with respect to $\bar{r}_{\min}$). Note that the support of $y(t)$ is contained in the time interval $0,NT+r_{\max}/c]$, and its spectrum is negligible outside the frequency interval $0,M\Delta+v_{\max}/\lambda]$.
Define the excess initial-range $d=\bar{d}-\bar{r}_{\min}\in[0,r_{\max}]$, the excess range-rate $v=\bar{v}-\bar{v}_{\min}\in[0,v_{\max}]$, and $\alpha=\bar{\alpha}\mathrm{e}^{-\mathrm{i}2\pi\frac{d}{\lambda}}$. Ignoring the variation of the target range in a symbol interval, then (11) can be expanded as
$y(t)=\sum_{n=0}^{N-1}y_{n}(t),$ (12)
where
$y_{n}(t)$ $=\alpha x_{n}\left(t-\frac{r\left(nT\right)}{c}\right)\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}t}$
$=\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}h_{n}(\tau,\nu)x_{n}\left(t-\tau\right)\mathrm{e}^{\mathrm{i}2\pi\nu(t-\tau)}d\tau d\nu$ (13)
and
$h_{n}(\tau,\nu)=\alpha\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(nT\right)}{c}}\delta\left(\tau-\frac{r\left(nT\right)}{c}\right)\delta\left(\nu-\frac{v}{\lambda}\right).$ (14)
Notice that $y_{n}(t)$ is the output of a doubly-selective channel with delay-Doppler spread function $h_{n}(\tau,\nu)$ when the input is $x_{n}(t)$; the spread function depends upon $n$, as the target range can change during the OTFS frame.
II-D OTFS demodulation
The radar receiver relies on an OTFS demodulator to obtain discrete samples of the received signal in the delay-Doppler domain. Such samples are then processed to estimate the target parameters as shown in Sec. IV.
First, the OTFS demodulator transforms $y(t)$ into the time-frequency domain via the Wigner Transform; in particular, the following measurements are obtained
$Y_{\rm TF}[n,m]=\int_{-\infty}^{\infty}y(t)g_{\rm rx}^{*}(t-nT)\mathrm{e}^{-\mathrm{i}2\pi m\Delta(t-nT)}dt,$ (15)
for $n=0,\ldots,N-1$ and $m=0,\ldots,M-1$, where $g_{\rm rx}(t)$ is the receive shaping filter with unit energy, effective temporal duration $T$ and effective spectral width $\Delta$. Denote by
$\gamma(\tau,\nu)=\int_{-\infty}^{\infty}g_{\rm tx}(\beta)g^{*}_{\rm rx}(\beta-\tau)\mathrm{e}^{-\mathrm{i}2\pi\nu(\beta-\tau)}d\beta$ (16)
the cross-ambiguity function of $g_{\rm tx}(t)$ and $g_{\rm rx}(t)$; then, the following proposition establishes a connection between $\{Y_{\rm TF}[n,m]\}$ and $\{X_{\rm TF}[n,m]\}$ that generalizes the previous result in [6, Theorem 1]. The proof is reported in Appendix A.
Proposition 1.
If the variation of the target range within each symbol interval can be ignored, then
$Y_{\rm TF}[n,m]=\sum_{n^{\prime}=0}^{N-1}\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n^{\prime},m^{\prime}]X_{\rm TF}[n^{\prime},m^{\prime}],$ (17)
for $n=0,\ldots,N-1$ and $m=0,\ldots,M-1$, where
$H_{n,m}[n^{\prime},m^{\prime}]=\alpha\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}nT}\mathrm{e}^{-\mathrm{i}2\pi m^{\prime}\Delta\frac{r\left(n^{\prime}T\right)}{c}}\\ \times\gamma\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c},(m-m^{\prime})\Delta-\frac{v}{\lambda}\right).$ (18)
Next, the OTFS demodulator transforms $Y_{\rm TF}[n,m]$ into the delay-Doppler domain via the Symplectic Finite Fourier Transform; in particular, we have
$Y_{\rm DD}[k,l]=\frac{1}{\sqrt{NM}}\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}Y_{\rm TF}[n,m]\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)},$ (19)
for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$.
III Delay–Doppler Input–Output Model
We establish here a relationship between $\{Y_{\rm DD}[k,l]\}$ and $\{X_{\rm DD}[k,l]\}$ for two representative shaping filters, namely ideal and rectangular filters. Ideal shaping filters lead to a simplified input–output relationship, and the resulting analysis provides useful insights into the impact of range migration and serves as a benchmark. In contrast, the more practical rectangular shaping filters yield a more involved relationship.
III-A Ideal shaping filters
Ideal shaping filters satisfies the bi-orthogonality condition [6, Sec. III]. In this case, the ambiguity function satisfies $\gamma(\tau,\nu)=1$ for $(\tau,\nu)\in[-\tau_{\max},\tau_{\max}]\times[-\nu_{\max},\nu_{\max}]$, and $\gamma(\tau,\nu)=0$ for $(\tau,\nu)\in[nT-\tau_{\max},nT+\tau_{\max}]\times[m\Delta-\nu_{\max},m\Delta+\nu_{\max}]$ for all integer pairs $(m,n)\neq(0,0)$. Consequently, (17) becomes
$Y_{\rm TF}[n,m]=H_{n,m}[n,m]X_{\rm TF}[n,m],$ (20)
for $n=1,\ldots,N-1$ and $m=0,\ldots,M-1$, where
$H_{n,m}[n,m]=\alpha\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}nT}\mathrm{e}^{-\mathrm{i}2\pi\frac{r\left(nT\right)}{c}m\Delta}.$ (21)
To proceed, denote by
$\mathcal{D}_{Q}(\nu)=\frac{1}{Q}\sum_{q=0}^{Q-1}\mathrm{e}^{-\mathrm{i}2\pi\nu q}=\frac{\mathrm{e}^{-\mathrm{i}2\pi\nu Q}-1}{Q\left(\mathrm{e}^{-\mathrm{i}2\pi\nu}-1\right)}$ (22)
the periodic $\mathrm{sinc}$ function (also called Dirichlet function) [^7] Recall that $\mathcal{D}_{Q}(\nu)$ is a periodic function of $\nu$ with period one and is zero at multiple of $1/Q$, except for $\nu=0,\pm 1,\pm 2,\ldots$ where it is one; also, $|\mathcal{D}_{Q}(\nu)|$ has a main-lobe centered at 0 of width $2/Q$ (between zero crossings) and Q2Q-2 sidelobes of width $1/Q$ (between zero crossings) in [0.5,0.5][-0.5,0.5]. The first sidelobe is approximately $13$ dB down from the main-lobe, and the subsequent sidelobes fall off at about $6$ dB per octave [39]. of degree $Q$. Then, the following result extends [6, Proposition 2] to the case of target range variations every $N/B$ symbol intervals. The proof is reported in Appendix B.
Proposition 2.
Assume that ideal shaping filters are employed. Under the blockwise constant-range approximation, we have
$Y_{\rm DD}[k,l]=\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}\!X_{\rm DD}[k^{\prime},l^{\prime}]\Phi[k-k^{\prime},l-l^{\prime}],$ (23)
for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$, where
$\Phi[k,l]$ $=\frac{1}{B}\sum_{b=0}^{B-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k}{N}-\frac{vT}{\lambda}\right)$
$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (24)
The following remarks and example provide further insight into Proposition 2.
Remark 1.
$\Phi[k,l]$ is periodic in $k$ and $l$ with period $N$ and $M$, respectively. This implies that (23) is a two-dimensional circular convolution, which can be rewritten as
$Y_{\rm DD}[k,l]$ $=\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\Phi\Big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\Big]$ (25a)
$=\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}\Big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\Big]\Phi[k^{\prime},l^{\prime}].$ (25b)
Here, $\Phi[k,l]$ is the convolution kernel relating the transmitted and received delay-Doppler symbols. Accordingly, $\alpha\Phi[k,l]$ represents the target response in the delay-Doppler domain, and $\Phi[k,l]$ can be interpreted as the associated point-spread function, specified by the target initial range and range-rate.
Remark 2.
Proposition 2 subsumes [6, Proposition 2] as a special case. To illustrate this point, first use (1) to write the range at time $bT_{B}$ and the range-rate as
$r\left(bT_{B}\right)$ $=(l_{b,{\rm int}}+l_{b,{\rm fra}})R_{\rm r},$ (26a)
$v$ $=(k_{\rm int}+k_{\rm fra})R_{\rm rr},$ (26b)
where $l_{b,{\rm int}}\in\{0,1,\ldots,M-1\}$, $k_{\rm int}\in\{0,1,\ldots,N-1\}$, and $l_{b,{\rm fra}},\,k_{\rm fra}\in(-0.5,0.5]$. Next, assume no range migration, whereby $l_{b,{\rm int}}=l_{{\rm int}}$ and $l_{b,{\rm fra}}=l_{{\rm fra}}$ for $b=0,\ldots,B-1$. Upon plugging (26) in (25b), we obtain
$Y_{\rm DD}[k,l]=\alpha\sum_{\bar{k}=-\lfloor N/2\rfloor}^{\lceil N/2\rceil-1}\sum_{\bar{l}=-\lfloor M/2\rfloor}^{\lceil M/2\rceil-1}\mathcal{D}_{N}\!\left(-\frac{\bar{k}+k_{\rm fra}}{N}\right)\\ \times\mathcal{D}_{M}\!\left(\frac{\bar{l}+l_{\rm fra}}{M}\right)X_{\rm DD}\Big[[k-k_{\rm int}+\bar{k}]_{N},[l-l_{\rm int}+\bar{l}]_{M}\Big],$ (27)
consistently with [6, Proposition 2]. In addition, if $k_{\rm fra}=l_{\rm fra}=0$, then (27) becomes
$Y_{\rm DD}[k,l]$ $=\alpha X_{\rm DD}\Big[[k-k_{\rm int}]_{N},[l-l_{{\rm int}}]_{M}\Big].$ (28)
Remark 3.
For large $N/B$ and $M$, the sum in (25) contains only a few dominant terms. Indeed, by substituting (26) into (24), we obtain
$\Phi[k,l]$ $=\frac{1}{B}\sum_{b=0}^{B-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k_{\rm int}-k_{\rm fra}}{B}\right)b}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k_{\rm int}-k_{\rm fra}}{N}\right)$
$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l_{b,{\rm int}}-l_{b,{\rm fra}}}{M}\right)\right).$ (29)
For large $Q$, $|\mathcal{D}_{Q}(\nu)|\approx 0$ for $\nu\notin[-D/Q,D/Q]$, with $0<2D+1\ll Q$, where the non-negative integer $D$ controls the approximation error [6]. Accordingly, $|\Phi[k,l]|\approx 0$ if $k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$.
Refer to caption    Refer to caption    Refer to caption
Figure 2: $|\Phi[k,l]|$ vs $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$, when $B=4$, $N=M=64$, $l_{0,{\rm int}}=42$, $k_{\rm int}=32$, and ideal shaping filters. Case 1 (left): $l_{1,{\rm int}}=l_{2,{\rm int}}=l_{3,{\rm int}}=42$ and $l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=k_{\rm fra}=0$. Case 2 (center): $l_{1,{\rm int}}=l_{2,{\rm int}}=l_{3,{\rm int}}=42$ and $l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=k_{\rm fra}=0.5$. Case 3 (right): $l_{1,{\rm int}}=43$, $l_{2,{\rm int}}=44$, $l_{3,{\rm int}}=45$, and $l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=k_{\rm fra}=0.5$.
Example 1.
To provide physical insight into the effect of range migration on the target point-spread function in (24), we consider a toy example with $B=4$, $N=M=64$, $l_{0,{\rm int}}=42$, and $k_{\rm int}=32$. Fig. 2 reports $|\Phi[k,l]|$ versus $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ for different target kinematics. In Case 1, we assume no range migration (i.e., $l_{1,{\rm int}}=l_{2,{\rm int}}=l_{3,{\rm int}}=42$), integer range (i.e., $l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=0$), and integer range-rate (i.e., $k_{\rm fra}=0$). It is seen that $|\Phi[k,l]|$ contains a single non-zero entry at $(k,l)=(32,42)$, as predicted by (28). In Case 2, we assume no range migration, fractional range with $l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=0.5$, and fractional range-rate with $k_{\rm fra}=0.5$. By comparing Cases 1 and 2, it is observed that fractional range and range-rate induce a spreading of $|\Phi[k,l]|$ along both the delay and Doppler dimensions, as predicted by (27). In Case 3, we assume that the target range varies four times over the OTFS frame, with $l_{1,{\rm int}}=43$, $l_{2,{\rm int}}=44$, $l_{3,{\rm int}}=45$ and $k_{\rm fra}=l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=0.5$. By comparing Cases 2 and 3, range migration produces an additional spreading of $|\Phi[k,l]|$, which constitutes the main novelty captured by Proposition 2. In all cases, most entries of $|\Phi[k,l]|$ remain significantly smaller than one, in agreement with Remark 3. Hence, the response of a high-speed target preserves a sparse structure, albeit with an expanded support.
III-A1 Vectorized model
The samples $\{Y_{\rm DD}[k,l]\}$ can be arranged into the vector $\bm{y}\in\mathbb{C}^{NM}$, where $(\bm{y})_{lN+k+1}=Y_{\rm DD}[k,l]$, for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$. From (25a), we obtain
$\bm{y}=\alpha\bm{\Phi}\bm{x},$ (30)
where $\bm{\Phi}\in\mathbb{C}^{NM\times NM}$ has entries $(\bm{\Phi})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\Phi\big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\big]$ and $\bm{x}\in\mathbb{C}^{NM}$ is defined by $(\bm{x})_{lN+k+1}=X_{\rm DD}[k,l]$, for $k,k^{\prime}=0,\ldots,N-1$ and $l,l^{\prime}=0,\ldots,M-1$. Alternatively, from (25b), we have
$\bm{y}=\alpha\bm{X}\bm{\phi},$ (31)
where $\bm{X}\in\mathbb{C}^{NM\times NM}$ has entries $(\bm{X})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=X_{\rm DD}\big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\big]$ and $\bm{\phi}\in\mathbb{C}^{NM}$ is defined by $(\bm{\phi})_{lN+k+1}=\Phi[k,l]$, for $k,k^{\prime}=0,\ldots,N-1$ and $l,l^{\prime}=0,\ldots,M-1$. Since the matrix $\bm{\Phi}$ and the vector $\bm{\phi}$ depend on the target initial-range and range-rate, we will denote them by $\bm{\Phi}(d,v)$ and $\bm{\phi}(d,v)$ when we want to make explicit this dependence.
III-B Rectangular shaping filters
Assume now $g_{\rm tx}(t)=g_{\rm rx}(t)=\sqrt{1/T}\,\mathrm{rect}(t/T)$. This implies that $\gamma(\tau,\nu)=0$ for $|\tau|\geq T$. Moreover, for large $M$, we have [6]
$\gamma(\tau,\nu)=\frac{T}{M}\sum_{p=0}^{M-1}g_{\rm tx}\left(\frac{pT}{M}\right)g^{*}_{\rm rx}\left(\frac{pT}{M}-\tau\right)\mathrm{e}^{-\mathrm{i}2\pi\nu\left(\frac{pT}{M}-\tau\right)}\\ =\begin{cases}\displaystyle\frac{1}{M}\sum_{p=0}^{\lceil M+\tau M/T\rceil-1}\!\!\mathrm{e}^{-\mathrm{i}2\pi\nu T\left(\frac{p}{M}-\frac{\tau}{T}\right)},&\tau\in(-T,0),\\ \displaystyle\frac{1}{M}\sum_{p=\lceil\tau M/T\rceil}^{M-1}\!\!\mathrm{e}^{-\mathrm{i}2\pi\nu T\left(\frac{p}{M}-\frac{\tau}{T}\right)},&\tau\in[0,T).\end{cases}$ (32)
Consequently, (17) simplifies as follows. For $n=1,\ldots,N-1$ and $m=0,\ldots,M-1$, we have
$Y_{\rm TF}[n,m]$ $=H_{n,m}[n,m]X_{\rm TF}[n,m]$
$\quad+\sum_{m^{\prime}=0,\,m^{\prime}\neq m}^{M-1}H_{n,m}[n,m^{\prime}]X_{\rm TF}[n,m^{\prime}]$
$\quad+\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n-1,m^{\prime}]X_{\rm TF}[n-1,m^{\prime}],$ (33)
where the second and third terms on the right-hand side correspond to ICI and ISI, respectively [6]. For $n=0$, instead, no ISI is present due to the guard interval between consecutive OTFS frames, and (17) reduces to
$Y_{\rm TF}[0,m]$ $=H_{0,m}[0,m]X_{\rm TF}[0,m]$
$\quad+\sum_{m^{\prime}=0,\,m^{\prime}\neq m}^{M-1}H_{0,m}[0,m^{\prime}]X_{\rm TF}[0,m^{\prime}],$ (34)
for $m=0,\ldots,M-1$.
The following result extends [6, Theorem 2] to the case where the target range is not an integer multiple of $R_{\rm r}$ and may vary every $N/B$ symbol intervals. The proof is provided in Appendix C.
Proposition 3.
Assume rectangular-shaped filters. Under the blockwise constant-range approximation, we have
$Y_{\rm DD}[k,l]=\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}\!X_{\rm DD}[k^{\prime},l^{\prime}]\Psi_{k^{\prime},l^{\prime}}[k-k^{\prime},l-l^{\prime}],$ (35)
for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$, where
$\Psi_{k^{\prime},l^{\prime}}[k,l]$ $=\frac{1}{B}\sum_{b=0}^{B-1}\Psi_{b,k^{\prime},l^{\prime}}[k,l],$ (36)
while
$\Psi_{b,k^{\prime},l^{\prime}}[k,l]=\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}}{M}}$
$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k}{N}-\frac{vT}{\lambda}\right)$
$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right),$ (37)
if $0\leq l^{\prime}\leq\big\lceil M-\frac{r\left(bT_{B}\right)M}{cT}\big\rceil-1$, and
$\Psi_{b,k^{\prime},l^{\prime}}[k,l]=\mathrm{e}^{-\mathrm{i}2\pi\frac{k^{\prime}}{N}}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}-M}{M}}$
$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\left[\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k}{N}-\frac{vT}{\lambda}\right)-\mathbbm{1}_{\{b=0\}}\frac{B}{N}\right]$
$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l+M}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right),$ (38)
if $\big\lceil M-\frac{r\left(bT_{B}\right)M}{cT}\big\rceil\leq l^{\prime}\leq M-1$.
The following remarks and example provide further insight into Proposition 3.
Remark 4.
The right-hand side of (35) is not a two-dimensional convolution, since $\Psi_{k^{\prime},l^{\prime}}[k,l]$ depends on $k^{\prime}$ and $l^{\prime}$ due to ISI and ICI. This is in sharp contrast with the ideal shaping filter case. For fixed $k^{\prime}$ and $l^{\prime}$, $\Psi_{k^{\prime},l^{\prime}}[k,l]$ acts as a location-dependent kernel relating the input symbol $X_{\rm DD}[k^{\prime},l^{\prime}]$ to the output symbols $Y_{\rm DD}[k,l]$, and is periodic in $k$ and $l$ with period $N$ and $M$, respectively. Accordingly, $\alpha\Psi_{k^{\prime},l^{\prime}}[k,l]$ represents the target response in the delay-Doppler domain to an input symbol located at $(k^{\prime},l^{\prime})$, and $\Psi_{k^{\prime},l^{\prime}}[k,l]$ can be interpreted as a location-dependent point-spread function, specified by the target initial range and range-rate.
Remark 5.
By using (26), the result in (35) can be recast as
$Y_{{\rm DD}}[k,l]=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{k^{\prime}=0}^{N-1}$
$\quad\times\Bigg[\sum_{l^{\prime}=0}^{M-l_{b,{\rm int}}-\lfloor l_{b,{\rm fra}}\rfloor-1}X_{\rm DD}[k^{\prime},l^{\prime}]\Psi_{b,k^{\prime},l^{\prime}}[k-k^{\prime},l-l^{\prime}]$
$\quad+\sum_{l^{\prime}=M-l_{b,{\rm int}}-\lfloor l_{b,{\rm fra}}\rfloor}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\Psi_{b,k^{\prime},l^{\prime}}[k-k^{\prime},l-l^{\prime}]\Bigg]$
$=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{\bar{k}=k_{\rm int}-k}^{N+k_{\rm int}-k-1}$
$\quad\times\Bigg[\sum_{\bar{l}_{b}=l_{b,{\rm int}}-l}^{M-l-1-\lfloor l_{b,{\rm fra}}\rfloor}\Psi_{b,k^{\prime},l^{\prime}}[k_{\rm int}-\bar{k},l_{b,{\rm int}}-\bar{l}_{b}]$
$\quad\times X_{\rm DD}\Big[[k-k_{\rm int}+\bar{k}]_{N},[l-l_{b,{\rm int}}+\bar{l}_{b}]_{M}\Big]$
$\quad+\sum_{\bar{l}_{b}=-l-\lfloor l_{b,{\rm fra}}\rfloor}^{l_{b,{\rm int}}-l-1}\Psi_{b,k^{\prime},l^{\prime}}[k_{\rm int}-\bar{k},l_{b,{\rm int}}-\bar{l}_{b}-M]$
$\quad\times X_{\rm DD}\Big[[k-k_{\rm int}+\bar{k}]_{N},[l-l_{b,{\rm int}}+\bar{l}_{b}]_{M}\Big]\Bigg]$
$=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{\bar{k}=-\lfloor N/2\rfloor}^{\lceil N/2\rceil-1}\sum_{\bar{l}_{b}=-l-\lfloor l_{b,{\rm fra}}\rfloor}^{M-l-1-\lfloor l_{b,{\rm fra}}\rfloor}$
$\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{N}\frac{l_{b,{\rm int}}+l_{b,{\rm fra}}}{M}}\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{N}\frac{l-l_{b,{\rm int}}+\bar{l}_{b}}{M}}$
$\quad\times\mathrm{e}^{\mathrm{i}2\pi\left(\frac{\bar{k}+k_{\rm fra}}{B}\right)b}A_{b,\bar{k},\bar{l}_{b}}[k,l]\mathcal{D}_{M}\!\left(\frac{\bar{l}_{b}+l_{b,{\rm fra}}}{M}\right)$
$\quad\times X_{\rm DD}\Big[[k-k_{\rm int}+\bar{k}]_{N},[l-l_{b,{\rm int}}+\bar{l}_{b}]_{M}\Big],$ (39)
where $A_{b,\bar{k},\bar{l}_{b}}[k,l]$ is equal to
$\mathcal{D}_{\frac{N}{B}}\!\left(-\frac{\bar{k}+k_{\rm fra}}{N}\right),$ (40)
if $l_{b,{\rm int}}-l\leq\bar{l}_{b}\leq M-l-1-\lfloor l_{b,{\rm fra}}\rfloor$, and to
$\left[\mathcal{D}_{\frac{N}{B}}\!\left(-\frac{\bar{k}+k_{\rm fra}}{N}\right)-\mathbbm{1}_{\{b=0\}}\frac{B}{N}\right]\mathrm{e}^{-\mathrm{i}2\pi\frac{[k-k_{\rm int}+\bar{k}]_{N}}{N}},$ (41)
if $l-\lfloor l_{b,{\rm fra}}\rfloor\leq\bar{l}_{b}\leq l_{b,{\rm int}}-l-1$. If $l_{b,{\rm int}}=l_{{\rm int}}$ and $l_{b,{\rm fra}}=0$ for $b=0,\ldots,B-1$, then (39) becomes
$Y_{{\rm DD}}[k,l]=\alpha\sum_{\bar{k}=-\lfloor N/2\rfloor}^{\lceil N/2\rceil-1}\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{N}\frac{l_{{\rm int}}}{M}}\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{N}\frac{l-l_{{\rm int}}}{M}}\\ \times A_{\bar{k}}[k,l]X_{\rm DD}\Big[[k-k_{\rm int}+\bar{k}]_{N},[l-l_{{\rm int}}]_{M}\Big],$ (42)
where $A_{\bar{k}}[k,l]$ is equal to
$\mathcal{D}_{N}\!\left(-\frac{\bar{k}+k_{\rm fra}}{N}\right),$ (43)
if $l_{{\rm int}}\leq l\leq M-1$, and to
$\left[\mathcal{D}_{N}\!\left(-\frac{\bar{k}+k_{\rm fra}}{N}\right)-\frac{1}{N}\right]\mathrm{e}^{-\mathrm{i}2\pi\frac{[k-k_{\rm int}+\bar{k}]_{N}}{N}},$ (44)
if $0\leq l\leq l_{{\rm int}}-1$, that is consistent with [6, Theorem 2].
Remark 6.
For large $N/B$ and $M$, the sum in (35) only contains few dominant terms; indeed, we have $|\Psi_{k^{\prime},l^{\prime}}[k,l]|\approx 0$ if $k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$, for $b=0,\ldots,B-1$. This result is similar to that obtained for ideal shaping filters in Remark 3.
Refer to caption    Refer to caption    Refer to caption
Refer to caption    Refer to caption    Refer to caption
Figure 3: $|\Psi_{k^{\prime},l^{\prime}}[k,l]|$ vs $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$, when $B=4$, $N=M=64$, and rectangular shaping filters are employed. Top and bottom rows are for $(k^{\prime},l^{\prime})=(0,0)$ and $(32,32)$, respectively. Cases 1 (left), 2 (center), and 3 (right) are defined as in Fig. 2.
Example 2.
Fig. 3 reports $|\Psi_{k^{\prime},l^{\prime}}[k,l]|$ versus $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ under the same parameter settings considered in Example 1, for the input locations $(l^{\prime},k^{\prime})=(0,0)$ and $(32,32)$. The resulting patterns show that the target responses exhibit similar overall support in the delay-Doppler domain, even though their shapes depend on the input location. For $(l^{\prime},k^{\prime})=(32,32)$, the Doppler profile is also affected by the term $\mathbbm{1}_{\{b=0\}}B/N$ in (38), which modifies the associated Dirichlet kernel due to the ISI induced by the rectangular shaping filters; this effect is not present for $(l^{\prime},k^{\prime})=(0,0)$.
III-B1 Vectorized model
The samples $\{Y_{\rm DD}[k,l]\}$ can again be arranged into the vector $\bm{y}\in\mathbb{C}^{NM}$. From (35), we obtain
$\bm{y}=\alpha\bm{\Psi}\bm{x},$ (45)
where $\bm{\Psi}\in\mathbb{C}^{NM\times NM}$ has entries $(\bm{\Psi})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\Psi_{k^{\prime},l^{\prime}}\big[k-k^{\prime},l-l^{\prime}\big]$. Note that the models in (30) and (45) share the same structure; however, the matrices $\bm{\Phi}$ and $\bm{\Psi}$ differ. We will denote $\bm{\Psi}$ by $\bm{\Psi}(d,v)$ when we want to make explicit its dependence on the target initial-range and range-rate.
The input-output relationship in (35), or equivalently (45), captures the combined effects of range migration, ICI, and ISI. We next introduce a more tractable approximated model that preserves the dominant delay-Doppler spreading effects induced by range migration while simplifying the dependence of the target response on the transmitted-symbol location.
III-B2 Approximated model
Notice first that (37) and (38) differ by the term
$e^{-\mathrm{i}2\pi\frac{k^{\prime}}{N}}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{d}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}-M}{M}}\\ \times\frac{B}{N}\mathcal{D}_{M}\!\left(-\left(\frac{l+M}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right)\mathbbm{1}_{\{b=0\}},$ (46)
whose magnitude tends to zero as $N/B$ increases. Upon neglecting the above term in (38) and ignoring the fractional part of the target range, then (39) is approximated as
$Y_{{\rm DD}}[k,l]\approx\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{NM}l}\Theta_{k^{\prime},l^{\prime}}[k,l]$
$\quad\times\frac{1}{B}\sum_{b=0}^{B-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k^{\prime}-k_{\rm int}-k_{\rm fra}}{B}\right)b}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k^{\prime}-k_{\rm int}-k_{\rm fra}}{N}\right)$
$\quad\times\mathcal{D}_{M}\!\left(-\frac{l^{\prime}-l_{b,{\rm int}}}{M}\right)X_{\rm DD}\Big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\Big],$ (47)
where
[…]
Based on (47), we can now write
$\bm{y}\approx\alpha\big(\bm{X}\odot\bm{\Theta}\odot\bm{\Lambda}(v)\big)\bm{\phi}(d,v),$ (49)
where $\bm{X}$ and $\bm{\phi}(d,v)$ are the same quantities in (31), while $\bm{\Theta}$ and $\bm{\Lambda}(v)$ are $NM\times NM$ matrices with $(\bm{\Theta})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\Theta_{k^{\prime},l^{\prime}}[k,l]$ and $\big(\bm{\Lambda}(v)\big)_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{NM}l}=\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda M}l}$, for $k,k^{\prime}=0,\ldots,N-1$ and $l,l^{\prime}=0,\ldots,M-1$. To proceed, notice that $\{\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda M}l}\}_{l=-\infty}^{\infty}$ is a complex exponential whose numerical frequency depends upon the target range-rate. Since $\frac{v_{\max}T}{\lambda}<1$ and the samples $\{\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda M}l}\}_{l=0}^{M-1}$ present an overall phase variation less than $2\pi$, we propose to approximate $\bm{\Lambda}(v)$ by $\bm{\Lambda}(v_{\max}/2)$. Hence, we finally obtain
$\bm{y}\approx\alpha\bm{\Xi}\bm{\phi}(d,v),$ (50)
where $\bm{\Xi}=\bm{X}\odot\bm{\Theta}\odot\bm{\Lambda}(v_{\max}/2)$. By comparing (31) with (50), it is seen that the symbol matrix $\bm{X}$ in (31) is now replaced in (49) by the modified symbol matrix $\bm{\Xi}$ to account for the presence of rectangular shaping filters.
To compare (45) and (50), we compute the corresponding normalized root mean square error (NRMSE), namely, $(\mathrm{E}\left[\|\bm{\Xi}\bm{\phi}(d,v)-\bm{\Psi}(d,v)\bm{x}\|^{2}\right]/\mathrm{E}\left[\|\bm{\Psi}(d,v)\bm{x}\|^{2}\right])^{1/2}$, where the statistical expectation is with respect to $\bm{x}$, $d$, and $v$. For the system setup considered in Sec. V, we found an NRMSE of about $0.07$ for any choice of the receiver design parameter $B$, thus confirming the goodness of the approximated model in (50).
IV Estimation of the target parameters
Upon accounting for the additive noise, after OTFS demodulation the received signal can be written as
$\bm{z}=\bm{y}+\bm{\omega}\in\mathbb{C}^{NM},$ (51)
where $\bm{\omega}$ is a circularly symmetric complex Gaussian vector with covariance matrix $\sigma_{\omega}^{2}\bm{I}_{NM}$. Based on $\bm{z}$ and the target echo model derived in Sec. III, the radar receiver is faced with the problem of estimating the unknown target parameters, namely $d$, $v$, and $\alpha$. The corresponding ML estimates are [40]
[…]
$\hat{\alpha}$ $=\frac{\bm{e}\big(\hat{d},\hat{v}\big)^{\mathsf{H}}\bm{z}}{\big\|\bm{e}\big(\hat{d},\hat{v}\big)\big\|^{2}},$ (52b)
where $\mathcal{S}=[0,r_{\max}]\times[0,v_{\max}]$, and
$\bm{e}(d,v)=\begin{cases}\bm{\Phi}(d,v)\bm{x},&\text{for ideal shaping filters},\\ \bm{\Psi}(d,v)\bm{x},&\text{for rectangular shaping filters}.\end{cases}$ (53)
In practice, the continuous set $\mathcal{S}$ is replaced by a uniform rectangular grid $\mathcal{G}$ containing $G_{\rm r}=\lfloor r_{\max}/\Sigma_{\rm r}\rfloor$ and $G_{\rm rr}=\lfloor v_{\max}/\Sigma_{\rm rr}\rfloor$ points along the initial-range and range-rate domains, respectively, where $\Sigma_{\rm r}\leq R_{\rm r}$ and $\Sigma_{\rm rr}\leq R_{\rm rr}$ are the corresponding point spacing.
A direct implementation of (52) entails a complexity of
$\mathcal{O}\big(G_{\rm r}G_{\rm rr}(MN)^{2}\big).$ (54)
In particular, the evaluation of (52) may become computationally demanding when a fine search grid is required. To alleviate this burden, we propose in Sec. IV-A a reduced-complexity two-step estimation procedure that exploits the sparse structure of the target echo. This procedure is further extended in Sec. IV-B to also handle multiple targets.
IV-A Two-step estimation
[…]
IV-A1 Coarse estimation of the initial-range and range-rate
Consider first ideal shaping filters. From (31), the received signal can be written as $\bm{z}=\bm{X}\bm{f}+\bm{\omega}$, where $\bm{f}=\alpha\bm{\phi}$ is the target response vector. For brevity, the dependence on $d$, $v$, and $B$ is omitted in this section. To proceed, $\bm{f}$ is partitioned into $M$ subvectors (blocks) of length $N$, namely, $\bm{f}=(\bm{f}_{0};\cdots;\bm{f}_{M-1})$, where $\bm{f}_{m}=\big(\alpha\Phi[0,m];\ldots;\alpha\Phi[N-1,m]\big)$. According to Remark 3, $\bm{f}$ is block-sparse for large $N/B$ and $M$. Next, the dimension of $\bm{z}$ is reduced by means of an $NK\times NM$ selection matrix $\bm{S}$ whose rows are distinct rows of $\bm{I}_{NM}$. The parameter $K\in\{1,\ldots,M\}$ controls the implementation complexity. The compressed measurement $\bm{z}_{s}=\bm{S}\bm{z}\in\mathbb{C}^{NK}$ can be expressed as
$\bm{z}_{s}=\bm{X}_{s}\bm{f}+\bm{\omega}_{s}=\sum_{m=0}^{M-1}\bm{X}_{s,m}\bm{f}_{m}+\bm{\omega}_{s},$ (55)
where $\bm{\omega}_{s}=\bm{S}\bm{\omega}$, $\bm{X}_{s}=(\bm{X}_{s,0}\,\cdots\,\bm{X}_{s,M-1})=\bm{S}\bm{X}$, and $\bm{X}_{s,0},\ldots,\bm{X}_{s,M-1}\in\mathbb{C}^{NK\times N}$.
Based on (55), an estimate of the target response vector, say $\tilde{\bm{f}}$, can be computed by resorting to the BOMP algorithm [31]. Let $\bm{z}^{(1)}_{s}=\bm{z}_{s}$ and $\mathcal{I}^{(1)}=\{0,\ldots,M-1\}$; then, at the $\ell$-th iteration, for $\ell\geq 1$, the index of the block in the dictionary matrix $\bm{X}_{s}$ best matched to $\bm{z}^{(\ell)}_{s}$ is first identified as follows
$i^{(\ell)}=\operatorname*{\arg\max}_{i\in\mathcal{I}^{(\ell)}}\Big\|\bm{X}_{s,i}^{\mathsf{H}}\bm{z}^{(\ell)}_{s}\Big\|^{2};$ (56)
then, an estimate of the sub-vectors $\bm{f}_{i^{(1)}},\ldots,\bm{f}_{i^{(\ell)}}$ corresponding to the currently selected blocks of $\bm{X}_{s}$ is computed by solving a least-square problem, namely,
$\big(\tilde{\bm{f}}_{i^{(1)}}\cdots\tilde{\bm{f}}_{i^{(\ell)}}\big)=\operatorname*{\arg\min}_{(\bm{f}_{i^{(1)}}\cdots\bm{f}_{i^{(\ell)}})\in\mathbb{C}^{N\times\ell}}\Big\|\bm{z}_{s}-\sum_{q=1}^{\ell}\bm{X}_{s,i^{(q)}}\bm{f}_{i^{(q)}}\Big\|^{2};$ (57)
finally, the residual vector and the search set are updated as $\bm{z}^{(\ell+1)}_{s}=\bm{z}^{(\ell)}_{s}-\sum_{q=1}^{\ell}\bm{X}_{s,i^{(q)}}\tilde{\bm{f}}_{i^{(q)}}$ and $\mathcal{I}^{(\ell+1)}=\mathcal{I}^{(\ell)}\setminus\{i^{(\ell)}\}$, respectively, and the procedure is stopped when $\ell=K$ or $\|\bm{z}^{(\ell+1)}_{s}\|^{2}<\epsilon NK\sigma_{\omega}^{2}$, with $\epsilon$ being a positive threshold. If the procedure ends in $\tilde{\ell}$ iterations, an estimate $\tilde{\bm{f}}$ of the target response vector is obtained from the indexes $i^{(1)},\ldots,i^{(\tilde{\ell})}$ and the sub-vectors $\tilde{\bm{f}}_{i^{(1)}},\ldots,\tilde{\bm{f}}_{i^{(\tilde{\ell})}}$. At this point, the entries of $\tilde{\bm{f}}$ are arranged into the sequence $\tilde{F}[k,l]=(\tilde{\bm{f}})_{lN+k+1}$, for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$, which provides an estimate of $\alpha\Phi[k,l]$; finally, coarse estimates of the target initial-range and range-rate are obtained from the location of the maximum element of $|\tilde{F}[k,l]|$.
When a rectangular shaping filter is employed, from (50), the received signal can be approximately expanded as $\bm{z}\approx\alpha\bm{\Xi}\bm{\phi}+\bm{\omega}$; accordingly, the above procedure can still be used after replacing $\bm{X}$ by the modified symbol matrix $\bm{\Xi}$.
IV-A2 Implementation complexity
We now discuss the implementation complexity of the proposed two-step estimator. In the first step, the BOMP algorithm is applied to the compressed model in (55). At the $\ell$-th iteration, the block-selection requires the computation of $\bm{X}_{s,i}^{\rm H}\bm{z}_{s}^{(\ell)}$ for all remaining blocks; since $\bm{X}_{s,i}\in\mathbb{C}^{NK\times N}$, this operation has complexity $\mathcal{O}((M-\ell+1)KN^{2})$. The following least-squares update involves the matrix obtained by concatenating the $\ell$ selected blocks; a standard dense least-squares implementation therefore requires $\mathcal{O}(K\ell^{2}N^{3}+\ell^{3}N^{3})$ operations. The residual updates have lower complexity and do not affect the dominant scaling. Hence, summing over at most $K$ iterations yields a complexity $\mathcal{O}(MK^{2}N^{2}+K^{4}N^{3})$ for the first step. In the second step, the metric in (52a) is evaluated over the local grid $\tilde{\mathcal{G}}$, with complexity $\mathcal{O}\big(\tilde{G}_{\rm r}\tilde{G}_{\rm rr}(MN)^{2}\big)$. Therefore, the overall complexity of the proposed two-step estimator is
$\mathcal{O}\!\left(MK^{2}N^{2}+K^{4}N^{3}+\tilde{G}_{\rm r}\tilde{G}_{\rm rr}(MN)^{2}\right).$ (58)
Since in practice $K\ll M$ and $\tilde{G}_{\rm r}\tilde{G}_{\rm rr}\ll G_{\rm r}G_{\rm rr}$, this complexity is substantially lower than that in (54).
IV-B Extension to multiple targets
Multiple targets can be handled through the iterative CLEAN algorithm [32, 33, 34, 35], which repeatedly estimates and removes the dominant target echo from the received signal. Let $\bm{r}^{(1)}=\bm{z}$ denote the initial residual. At the $p$-th iteration, the proposed two-step estimator is applied to $\bm{r}^{(p)}$ to extract the dominant remaining target and estimate its initial-range $\hat{d}^{(p)}$, range-rate $\hat{v}^{(p)}$, and amplitude $\hat{\alpha}^{(p)}$. The corresponding echo is then reconstructed as $\hat{\bm{y}}^{(p)}=\hat{\alpha}^{(p)}\bm{e}\big(\hat{d}^{(p)},\hat{v}^{(p)}\big)$, and the residual is updated as $\bm{r}^{(p+1)}=\bm{r}^{(p)}-\hat{\bm{y}}^{(p)}$ to mitigate the interference caused by previously extracted targets in the next iteration $p+1$. If the number of targets is known, say $P$, the procedure is repeated exactly $P$ times. Otherwise, the iterations are stopped when $\|\bm{r}^{(p+1)}\|^{2}$ falls below a given threshold or when a maximum number of iterations is reached.
Since each iteration requires one execution of the proposed two-step estimator, with complexity given in (58), together with a residual update of complexity $\mathcal{O}(MN)$, the overall complexity scales linearly with the number of extracted targets. If $P$ targets are processed, the resulting complexity is
$\mathcal{O}\big(P\big(MK^{2}N^{2}+K^{4}N^{3}+\tilde{G}_{\rm r}\tilde{G}_{\rm rr}(MN)^{2}\big)\big).$ (59)
V Numerical analysis
To assess the system performance, we consider the parameters in Table I. For these values, the maximum variation of the target range during one OTFS frame is $\bar{v}_{\rm max}NT\approx 136$ m, which corresponds to about $3.5$ times the range resolution $R_{\rm r}$. The initial-range $\bar{d}$ and range-rate $\bar{v}$ are randomly generated within the intervals $\bar{r}_{\rm min},\,\bar{r}_{\rm max}]$ and $\bar{v}_{\rm min},\,\bar{v}_{\rm max}]$, respectively, while a Swerling I fluctuation model is adopted for the amplitude $\alpha$. The information symbols $\{X_{\rm DD}[k,l]\}$ are independently drawn from a unit-energy $4$-QAM constellation.
The entries of the observed vector $\bm{z}$ are generated according to (12), (15), and (19). The radar receiver, instead, is designed according to the model developed in Secs. III-A and III-B for ideal and rectangular shaping filters, respectively. Unless otherwise stated, we set $B=16$. In the implementation of the proposed two-step estimator, we use $\Sigma_{\rm r}=R_{\rm r}/100$, $\Sigma_{\rm rr}=R_{\rm rr}/100$, $\tilde{G}_{\rm r}=401$, $\tilde{G}_{\rm rr}=401$, $K=4P$, and $\epsilon=1$.
For future reference, we define the target SNR as
$\mathrm{SNR}=\mathrm{E}[|\alpha|^{2}]NM/\sigma_{\omega}^{2}.$ (60)
Also, let $d$, $v$, and $\alpha$ denote the true values of the initial-range, range-rate, and amplitude of the target, and let $\hat{d}$, $\hat{v}$, and $\hat{\alpha}$ be the corresponding estimates. We define the root mean square error (RMSE) of the initial-range and range-rate estimates, and the NRMSE of the amplitude estimate, as $\mathrm{RMSE}_{\rm r}=(\mathrm{E}[|d-\hat{d}|^{2}])^{1/2}$, $\mathrm{RMSE}_{\rm rr}=(\mathrm{E}[|v-\hat{v}|^{2}])^{1/2}$, and $\mathrm{NRMSE}_{\rm\alpha}=(\mathrm{E}[|\alpha-\hat{\alpha}|^{2}]/\mathrm{E}[|\alpha|^{2}])^{1/2}$, respectively. In the following, we first examine the single-target case and then the multi-target scenario.
Table I: System parameters
$c/\lambda$ Carrier frequency 4 GHz
$\Delta$ Subcarrier spacing 15 kHz
$M$ Number of subcarriers 512
NN Number of symbol intervals 128
$R_{\rm r}$ Range resolution $\approx$ 39 m
$R_{\rm rr}$ Range-rate resolution $\approx$ 8.8 m/s
$\bar{r}_{\rm min},\,\bar{r}_{\rm max}]$ Inspected initial-range interval [20, 39] km
$\bar{v}_{\rm min},\,\bar{v}_{\rm max}]$ Inspected range-rate interval [15, 16] km/s
V-A Single target
Refer to caption Refer to caption Refer to caption
Figure 4: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $B=1,2,4,8,16,32,64,128$ for $\mathrm{SNR}=10,20,30$ dB, when a high- or low-speed target is present and ideal shaping filters are employed. Both the ML and the proposed two-step estimator are considered for comparison.
Refer to caption Refer to caption Refer to caption
Figure 5: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $\mathrm{SNR}$ for $B=4,8,16$, when a high-speed target is present, and the proposed two-step estimator is employed. Both ideal and rectangular shaping filters are considered.
Refer to caption Refer to caption Refer to caption
Figure 6: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $B=1,2,4,8,16,32,64,128$ for different values NN and $\Delta$, when a high-speed target is present, rectangular shaping filters are employed, $\mathrm{SNR}=30$ dB, and the proposed two-step estimator is employed.
Refer to caption Refer to caption Refer to caption
Figure 7: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $\bar{v}_{\rm min}$, when rectangular filters are used, $\mathrm{SNR}=15$ dB, and $\bar{v}_{\rm max}=\bar{v}_{\rm min}+1$ km/s. The proposed two-step estimator is implemented for $B=1,2,4,8,16$. For comparison, the threshold-based estimator in [17] and the SOMP-based estimator in [24] are also included.
We first examine the system performance with ideal shaping filters. We then consider rectangular shaping filters.
V-A1 Ideal shaping filters
Fig. 4 presents the metrics $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\rm\alpha}$ as functions of $B$ for $\mathrm{SNR}=10,20,30$ dB. Observe that the parameters of a high-speed target cannot be accurately recovered when range variation during the OTFS frame is neglected at the system design (i.e., $B=1$). As $B$ increases from 1 to 16, estimation accuracy improves significantly, while further increases in $B$ yield diminishing returns. This behavior is governed by the ratio between the maximum range variation during one block of $N/B$ symbols and the range resolution, namely
$\frac{\bar{v}_{\max}NT/B}{R_{\rm r}}=\frac{\bar{v}_{\max}MN}{cB},$ (61)
which should be sufficiently smaller than one for accurate modeling. For the considered parameters, this ratio equals approximately $3.5/B$. Hence, for $B=1$, the blockwise constant-range approximation is inaccurate, whereas for $B=16$ the ratio reduces to about $0.22$, corresponding to a maximum range variation over $N/B$ symbols of approximately $8.5$ m. Such variation can be neglected with little performance loss.
[…]
Finally, note that the proposed two-step estimator achieves performance close to that of the ML estimator across all scenarios considered in Fig. 4, at a much lower implementation complexity, underscoring its effectiveness and practical utility.
V-A2 Rectangular shaping filters
Fig. 5 shows the behavior of $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\rm\alpha}$ as functions of $\mathrm{SNR}$ when the proposed two-step estimator is employed and $B=4,8,16$. For comparison, results obtained with ideal shaping filters are also shown. Notably, the more practical rectangular filters achieve performance close to that of the ideal filters.
Fig. 6 investigates the influence of the OTFS parameters $\Delta$ and $N$ on the estimation performance when the proposed two-step estimator is employed and $\mathrm{SNR}=30$ dB. The results show that the initial-range estimation accuracy improves as the subcarrier spacing $\Delta$ increases, since the occupied bandwidth $M\Delta$ becomes larger, whereas it is only marginally affected by variations of $N$ when SNR is kept constant. In contrast, the range-rate estimation accuracy benefits from increasing $N$ and reducing $\Delta$, since both choices enlarge the coherent observation time $NT$ (recall that $T=1/\Delta$). As for the impact of the parameter $B$, the ratio introduced in (61) remains the key quantity. Note that this ratio is independent of $\Delta$ and scales linearly with $N$. Therefore, the pairs $(N,\Delta)=(128,15\penalty 10000\ \text{kHz})$ and $(128,30\penalty 10000\ \text{kHz})$, represented by solid lines, exhibit a similar validity region for the blockwise constant-range approximation, whereas the remaining pairs with $N=256$, represented by dashed lines, require approximately twice the value of $B$ to achieve a comparable level of approximation accuracy.
Finally, we evaluate the estimation performance of the proposed two-step estimator across varying mobility scenarios; to this end, we set $\bar{v}_{\rm max}=\bar{v}_{\rm min}+1$ km/s and vary $\bar{v}_{\rm min}$ from 0 to 15 km/s. For comparison, we also include the threshold-based estimator in [17] and the SOMP-based estimator in [24], both of which are designed under the assumption of a constant integer range within the OTFS frame. To ensure a fair comparison, the input delay-Doppler sequence $X_{{\rm DD}}[k,l]$ contains a single non-zero entry located at $(k,l)=(64,256)$, whereas all remaining entries are set to zero to create a guard region. [^8] In this setup, the complexity of the proposed two-step estimator simplifies to $\mathcal{O}\big(MK^{2}N^{2}+K^{4}N^{3}+\tilde{G}_{\text{r}}\tilde{G}_{\text{rr}}MN\big)$ because the vector $\bm{x}$ contains a single non-zero entry at a known location and the matrix-vector multiplication in (53) becomes trivial during the second step. In contrast, the benchmarks from [17] and [24] both have complexity $\mathcal{O}(MN)$. Fig. 7 plots the metrics $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\alpha}$ versus $\bar{v}_{\rm min}$ for $\mathrm{SNR}=15$ dB. The performance of the estimators in [17] and [24] deteriorates rapidly when $\bar{v}_{\rm min}$ exceeds 1 km/s. In this regime, the maximum range variation during the OTFS frame, given by $\bar{v}_{\rm max}NT$, becomes comparable to or exceeds the range resolution $R_{\rm r}$, resulting in a mismatch between the assumed and actual echo models. In contrast, the proposed two-step estimator remains robust over a significantly wider mobility range by using larger values of $B$, which enable a finer blockwise constant-range approximation in (8). In particular, $B=16$ provides stable performance over all considered values of $\bar{v}_{\rm min}$, consistently with the trends already observed in Figs. 4 and 6. Even when $\bar{v}_{\rm min}\rightarrow 0$, the proposed estimator still outperforms the existing methods due to the second refinement step, which improves the estimation accuracy at the cost of higher computational complexity.
V-B Multiple targets
Refer to caption Refer to caption Refer to caption
Figure 8: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $\mathrm{SNR}_{2}$, when rectangular filters are used, $P=3$ high-speed targets are present, $B=16$, and $\mathrm{SNR}_{1}/\mathrm{SNR}_{2}=\mathrm{SNR}_{2}/\mathrm{SNR}_{3}=10$. The performance of the proposed CLEAN-based estimator and the corresponding single-target performance bound are reported.
Refer to caption Refer to caption Refer to caption
Figure 9: $|\tilde{F}[k,l]|/\sigma_{\omega}$ vs $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ at the first (left), second (center), and third (right) iteration of the CLEAN procedure, when rectangular filters are employed, $P=3$ high-speed targets are present, $B=16$, $\mathrm{SNR}_{1}=30$ dB, $\mathrm{SNR}_{2}=20$ dB, and $\mathrm{SNR}_{3}=10$ dB.
Refer to caption Refer to caption Refer to caption
Figure 10: Average $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus range-rate separation, when rectangular filters are used, $P=2$ high-speed targets with the same initial-range are present, $B=16$, and $\mathrm{SNR}_{1}=\mathrm{SNR}_{2}=30$ dB. The performance of the proposed CLEAN-based estimator and the corresponding single-target performance bound are reported.
We use rectangular-shaped filters here and consider $P=3$ targets. Let $\mathrm{SNR}_{p}$ be the SNR of Target $p$ for $p=1,\ldots,P$. For each target, Fig. 8 reports the metrics $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\rm\alpha}$ versus $\mathrm{SNR}_{2}$, when $B=16$, $\mathrm{SNR}_{1}/\mathrm{SNR}_{2}=\mathrm{SNR}_{2}/\mathrm{SNR}_{3}=10$, and the proposed CLEAN-based estimator is employed. For comparison, we also plot the achievable performance when only a single target is present and the ML estimator in (52) is employed: we refer to it as the single-target bound. The proposed procedure effectively estimates the parameters of each target, achieving performance close to the single-target bound. For a single instance and $\mathrm{SNR}_{2}=20$ dB, Fig. 9 shows $|\tilde{F}[k,l]|/\sigma_{\omega}$ versus $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ in each iteration of the CLEAN procedure, where $\tilde{F}[k,l]$ is the sequence recovered by the BOMP algorithm. It is seen that the target echoes are iteratively recovered and subtracted out from the observed data.
Finally, we assess the ability of the proposed CLEAN-based estimator to resolve two targets having the same initial-range and SNR but different range-rates. Fig. 10 reports $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\alpha}$ (averaged over the two targets) versus the range-rate separation. When the range-rate separation is below the nominal resolution $R_{\rm rr}$, the two targets are difficult to distinguish, and the estimation errors increase significantly. As the separation increases beyond $R_{\rm rr}$, the two targets become reliably separable, and the performance approaches the single-target bound. This confirms that $R_{\rm rr}$ provides a meaningful reference for range-rate discrimination.
VI Conclusions
In this work, we have investigated the use of OTFS communication signals for estimating the parameters of space targets experiencing range migration, as may arise in LEO satellite applications. Our main technical contribution is the derivation of a novel echo model that allows the target range to vary $B$ times within one OTFS frame. In particular, we have shown that range migration induces a spreading of the target response while preserving its sparse structure in the delay-Doppler domain. We have also developed a two-step estimator for the target initial-range, range-rate, and amplitude, achieving performance close to that of the maximum likelihood estimator at substantially lower complexity. When combined with the CLEAN algorithm, the proposed procedure can handle multiple targets. A significant finding is that the achievable estimation accuracy for high-speed targets is comparable to that of low-speed targets when range migration is properly accounted for in the receiver design.
Future studies should consider that the radar receiver may not have perfect knowledge of the emitted communication symbols. In such cases, dedicated radar symbols may be considered, and the resource allocation between sensing and communication tasks should be investigated. Furthermore, using multiple coordinated LEO satellites for signal transmission and reception may allow for estimating the initial position and velocity vectors of the target rather than only its initial-range and range-rate. Finally, the use of multiple OTFS frames of different sizes may provide additional degrees of freedom and help resolve possible initial-range and range-rate ambiguities.
Appendix A Proof of Proposition 1
Upon plugging (12) into (15), we have
$Y_{\rm TF}[n,m]=\sum_{n^{\prime}=0}^{N-1}\sum_{m^{\prime}=0}^{M-1}\alpha X_{\rm TF}[n^{\prime},m^{\prime}]$
$\quad\times\int_{-\infty}^{\infty}g_{\rm tx}\!\left(t-n^{\prime}T-\frac{r\left(n^{\prime}T\right)}{c}\right)g^{*}_{\rm rx}\!\left(t-nT\right)$
$\quad\times\mathrm{e}^{\mathrm{i}2\pi m^{\prime}\Delta\left(t-n^{\prime}T-\frac{r\left(n^{\prime}T\right)}{c}\right)}\mathrm{e}^{-\mathrm{i}2\pi m\Delta\left(t-nT\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}t}dt$
$=\sum_{n^{\prime}=0}^{N-1}\sum_{m^{\prime}=0}^{M-1}X_{\rm TF}[n^{\prime},m^{\prime}]$
$\quad\times\alpha\mathrm{e}^{\mathrm{i}2\pi m\Delta\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c}\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\left(n^{\prime}T+\frac{r\left(n^{\prime}T\right)}{c}\right)}$
$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c}\right)}$
$\quad\times\int_{-\infty}^{\infty}g_{\rm tx}\!\left(\beta\right)g^{*}_{\rm rx}\!\left(\beta-(n-n^{\prime})T+\frac{r\left(n^{\prime}T\right)}{c}\right)$
$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left(\beta-(n-n^{\prime})T+\frac{r\left(n^{\prime}T\right)}{c}\right)}d\beta$
$=\sum_{n^{\prime}=0}^{N-1}\sum_{m^{\prime}=0}^{M-1}X_{\rm TF}[n^{\prime},m^{\prime}]\alpha\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}nT}\mathrm{e}^{-\mathrm{i}2\pi m^{\prime}\Delta\frac{r\left(n^{\prime}T\right)}{c}}$
$\quad\times\gamma\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c},(m-m^{\prime})\Delta-\frac{v}{\lambda}\right),$
which completes the proof.
Appendix B Proof of Proposition 2
After exploiting (3) and (20), Eq. (19) becomes
$Y_{\rm DD}[k,l]=\frac{1}{NM}\!\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}\!X_{\rm DD}[k^{\prime},l^{\prime}]\\ \times\underbrace{\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}H_{n,m}[n,m]\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}\right)n}\mathrm{e}^{\mathrm{i}2\pi\left(\frac{l-l^{\prime}}{M}\right)m}}_{(\star)}.$ (62)
Finally, notice that we have
$(\star)$ $=M\alpha\sum_{n=0}^{N-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)n}$
$\quad\times\left[\frac{1}{M}\sum_{m=0}^{M-1}\mathrm{e}^{\mathrm{i}2\pi\left(\frac{l-l^{\prime}}{M}-\frac{r\left(nT\right)\Delta}{c}\right)m}\right]$
$=M\alpha\sum_{n=0}^{N-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)n}$
$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(nT\right)\Delta}{c}\right)\right)$
$=NM\alpha\sum_{b=0}^{B-1}\frac{1}{B}\left[\frac{B}{N}\sum_{n=bN/B}^{(b+1)N/B-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)n}\right]$
$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right)$ (63)
$=NM\alpha\Phi[k-k^{\prime},l-l^{\prime}],$
where (63) follows from (8).
Appendix C Proof of Proposition 3
Upon plugging (33) and (34) in (19), we have
$Y_{\rm DD}[k,l]$ $=\frac{1}{\sqrt{NM}}\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)}$
$\quad\times\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n,m^{\prime}]X_{\rm TF}[n,m^{\prime}]$
$\quad+\frac{1}{\sqrt{NM}}\sum_{n=1}^{N-1}\sum_{m=0}^{M-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)}$
$\quad\times\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n-1,m^{\prime}]X_{\rm TF}[n-1,m^{\prime}]$
$=Y_{\rm DD}^{\rm ici}[k,l]+Y_{\rm DD}^{\rm isi}[k,l].$ (64)
As to the term $Y_{\rm DD}^{\rm ici}[k,l]$ in (64), we have:
$Y_{\rm DD}^{\rm ici}[k,l]$ $=\frac{1}{NM}\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)}\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n,m^{\prime}]$
$\quad\times\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{\mathrm{i}2\pi\left(\frac{nk^{\prime}}{N}-\frac{m^{\prime}l^{\prime}}{M}\right)}$
$=\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\Pi_{k,l}^{\rm ici}[k^{\prime},l^{\prime}],$ (65)
where
$\Pi_{k,l}^{\rm ici}[k^{\prime},l^{\prime}]$ $=\frac{1}{NM}\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n,m^{\prime}]$
$\quad\times\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}\right)}e^{j2\pi\left(\frac{ml-m^{\prime}l^{\prime}}{M}\right)}.$ (66)
$\Pi_{k,l}^{\rm ici}[k^{\prime},l^{\prime}]=\frac{\alpha}{NM^{2}}\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}\sum_{m^{\prime}=0}^{M-1}\left[\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}nT}\mathrm{e}^{-\mathrm{i}2\pi m^{\prime}\Delta\frac{r\left(nT\right)}{c}}\sum_{p=0}^{\big\lceil M-\frac{r\left(nT\right)M}{cT}\big\rceil-1}\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left(\frac{pT}{M}+\frac{r\left(nT\right)}{c}\right)}\right]\mathrm{e}^{\mathrm{i}2\pi n\left(\frac{k^{\prime}-k}{N}\right)}\mathrm{e}^{\mathrm{i}2\pi\left(\frac{ml-m^{\prime}l^{\prime}}{M}\right)}$
$=\frac{\alpha}{N}\sum_{n=0}^{N-1}\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(nT\right)}{c}}\sum_{p=0}^{\big\lceil M-\frac{r\left(nT\right)M}{cT}\big\rceil-1}\left[\frac{1}{M}\sum_{m=0}^{M-1}\mathrm{e}^{-\mathrm{i}2\pi\frac{m}{M}\left(p+\frac{r\left(nT\right)}{c}\frac{M}{T}-l\right)}\right]\left[\frac{1}{M}\sum_{m^{\prime}=0}^{M-1}\mathrm{e}^{\mathrm{i}2\pi\frac{m^{\prime}}{M}(p-l^{\prime})}\right]\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{p}{M}}$
$=\sum_{b=0}^{B-1}\frac{\alpha}{N}\sum_{n=bN/B}^{(b+1)N/B-1}\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(nT\right)}{c}}\sum_{p=0}^{\big\lceil M-\frac{r\left(nT\right)M}{cT}\big\rceil-1}\mathcal{D}_{M}\!\left(\frac{p+\frac{r\left(nT\right)}{c}\frac{M}{T}-l}{M}\right)\delta\big[[p-l^{\prime}]_{M}\big]\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{p}{M}}$
$=\frac{\alpha}{B}\sum_{b=0}^{B-1}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\mathcal{D}_{\frac{N}{B}}\!\left(\tfrac{k-k^{\prime}}{N}-\tfrac{vT}{\lambda}\right)\sum_{p=0}^{\big\lceil M-\frac{r\left(bT_{B}\right)M}{cT}\big\rceil-1}\mathcal{D}_{M}\!\left(\tfrac{p+\frac{r\left(bT_{B}\right)}{c}\frac{M}{T}-l}{M}\right)\delta\big[[p-l^{\prime}]_{M}\big]\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{p}{M}}.$ (67)
Upon expanding $\Pi_{k,l}^{\rm ici}[k^{\prime},l^{\prime}]$ as shown in (67) at the top of the next page, we obtain
$Y_{{\rm DD}}^{\rm ici}[k,l]$ $=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{\big\lceil M-\frac{r\left(bT_{B}\right)M}{cT}\big\rceil-1}X_{\rm DD}[k^{\prime},l^{\prime}]$
$\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}}{M}}$
$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)$
$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (68)
As to the term $Y_{\rm DD}^{\rm isi}[k,l]$ in (64), we have:
$Y_{\rm DD}^{\rm isi}[k,l]$ $=\frac{1}{NM}\sum_{n=1}^{N-1}\sum_{m=0}^{M-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)}\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n-1,m^{\prime}]$
$\quad\times\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{\mathrm{i}2\pi\left(\frac{(n-1)k^{\prime}}{N}-\frac{m^{\prime}l^{\prime}}{M}\right)}$
$=\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\Pi_{k,l}^{\rm isi}[k^{\prime},l^{\prime}]\mathrm{e}^{-\mathrm{i}2\pi\frac{k^{\prime}}{N}},$ (69)
where
$\Pi_{k,l}^{\rm isi}[k^{\prime},l^{\prime}]$ $=\frac{1}{NM}\sum_{n=1}^{N-1}\sum_{m=0}^{M-1}\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n-1,m^{\prime}]$
$\quad\times\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}\right)}e^{j2\pi\left(\frac{ml-m^{\prime}l^{\prime}}{M}\right)}.$ (70)
$\Pi_{k,l}^{\rm isi}[k^{\prime},l^{\prime}]=\frac{\alpha}{NM^{2}}\!\sum_{n=1}^{N-1}\sum_{m=0}^{M-1}\sum_{m^{\prime}=0}^{M-1}\!\left[\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}nT}\mathrm{e}^{-\mathrm{i}2\pi m^{\prime}\Delta\frac{r\left((n-1)T\right)}{c}}\!\!\!\!\sum_{p=\left\lceil M-\frac{r\left((n-1)T\right)M}{cT}\right\rceil}^{M-1}\!\!\!\!\!\!\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left(\frac{pT}{M}-T+\frac{r\left((n-1)T\right)}{c}\right)}\!\right]\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}\right)}\!\mathrm{e}^{\mathrm{i}2\pi\left(\frac{ml-m^{\prime}l^{\prime}}{M}\right)}$
$=\frac{\alpha}{N}\sum_{n=1}^{N-1}\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left((n-1)T\right)}{c}}\!\!\sum_{p=\left\lceil M-\frac{r\left((n-1)T\right)}{c}\frac{M}{T}\right\rceil}^{M-1}\!\!\left[\frac{1}{M}\sum_{m=0}^{M-1}\mathrm{e}^{-\mathrm{i}2\pi\frac{m}{M}\left(p-M+\frac{r\left((n-1)T\right)}{c}\frac{M}{T}-l\right)}\right]\left[\frac{1}{M}\sum_{m^{\prime}=0}^{M-1}\mathrm{e}^{\mathrm{i}2\pi\frac{m^{\prime}}{M}(p-M-l^{\prime})}\right]\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{p-M}{M}}$
$=\sum_{b=0}^{B-1}\frac{\alpha}{N}\sum_{\begin{subarray}{c}n=bN/B\\ n\neq 0\end{subarray}}^{(b+1)N/B-1}\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left((n-1)T\right)}{c}}\sum_{p=\left\lceil M-\frac{r\left((n-1)T\right)}{c}\frac{M}{T}\right\rceil}^{M-1}\mathcal{D}_{M}\!\left(\frac{p-M+\frac{r\left((n-1)T\right)}{c}\frac{M}{T}-l}{M}\right)\delta\big[[p-l^{\prime}]_{M}\big]\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{p-M}{M}}$
$=\frac{\alpha}{B}\sum_{b=0}^{B-1}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\left[\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)-\mathbbm{1}_{\{b=0\}}\frac{B}{N}\right]$
$\qquad\times\sum_{p=\left\lceil M-\frac{r\left(bT_{B}\right)}{c}\frac{M}{T}\right\rceil}^{M-1}\mathcal{D}_{M}\!\left(\frac{p-M+\frac{r\left(bT_{B}\right)}{c}\frac{M}{T}-l}{M}\right)\delta\big[[p-M-l^{\prime}]_{M}\big]\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{p-M}{M}}.$ (71)
Upon expanding $\Pi_{k,l}^{\rm isi}[k^{\prime},l^{\prime}]$ as shown in (71) at the top of the next page, we obtain
$Y_{{\rm DD}}^{\rm isi}[k,l]$ $=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=\big\lceil M-\frac{r\left(bT_{B}\right)M}{cT}\big\rceil}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{-\mathrm{i}2\pi\frac{k^{\prime}}{N}}$
$\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}-M}{M}}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}$
$\quad\times\left[\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)-\mathbbm{1}_{\{b=0\}}\frac{B}{N}\right]$
$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}+M}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (72)
Finally, (35) follows from (68) and (72).